ARevisedVersionofms BillTuthill ComputingServices UniversityofCalifornia Berkeley,CA94720 Themsmacroshavebeenslightlyrevisedandrearranged.Becauseoftherearrangement,thenew macroscanbereadbythecomputerinabouthalfthetimerequiredbythepreviousversionofms.This meansthatoutputwillbegintoappearbetweentensecondsandseveralminutesmorequickly,depending onthesystemload.Onlongfiles,however,thesavingsintotaltimearenotsubstantial.Theoldversion ofmsisstillavailableasmos. Severalbugsinmshavebeenfixed,includingabadproblemwiththe.1Cmacro,minordifficulties withboxedtext,abreakinducedby.EQbeforeinitialization,thefailuretosettabstopsindisplays,and severalbothersomeerrorsintherefermacros.MacrosusedonlyatBellLaboratorieshavebeenremoved. Thereareafewextensionstopreviousmsmacros,andanumberofnewmacros,butallthedocumented msmacrosstillworkexactlyastheydidbefore,andhavethesamenamesasbefore.Outputproduced withmsshouldlooklikeoutputproducedwithmos. Oneimportantnewfeatureisautomaticallynumberedfootnotes.Footnotenumbersareprintedby meansofapre-definedstring(\∗∗),whichyouinvokeseparatelyfrom.FSand.FE.Eachtimeitisused, thisstringincreasesthefootnotenumberbyone,whetherornotyouuse.FSand.FEinyourtext.Footnotenumberswillbesuperscriptedonthephototypesetterandondaisy-wheelterminals,butonlow-resolutiondevices(suchasthelprandacrt),theywillbebracketed.Ifyouuse\∗∗toindicatenumbered footnotes,thenthe.FSmacrowillautomaticallyincludethefootnotenumberatthebottomofthepage. Thisfootnote,forexample,wasproducedasfollows:1 Thisfootnote,forexample,wasproducedasfollows:\∗∗ .FS ... .FE Ifyouareusing\∗∗tonumberfootnotes,butwantaparticularfootnotetobemarkedwithanasteriskor adagger,thengivethatmarkasthefirstargumentto.FS:† thengivethatmarkasthefirstargumentto.FS:\(dg .FS\(dg ... .FE Footnotenumberingwillbetemporarilysuspended,becausethe\∗∗stringisnotused.Insteadofadagger,youcoulduseanasterisk*ordoubledagger‡,representedas\(dd. AnothernewfeatureisamacroforprintingthesesaccordingtoBerkeleystandards.Thismacrois called.TM,whichstandsforthesismode.(Itismuchlikethe.thmacroinme.)Itwillputpagenumbersintheupperright-handcorner;numberthefirstpage;suppressthedate;anddoublespaceeverything exceptquotes,displays,andkeeps.Useitatthetopofeachfilemakingupyourthesis.Calling.TM definesthe.CTmacroforchaptertitles,whichskipstoanewpageandmovesthepagenumbertothecenterfooter.The.P1(Pone)macrocanbeusedevenwithoutthesismodetoprinttheheaderonpage1, 1Ifyouneverusethe‘‘\∗∗’’string,nofootnotenumberswillappearanywhereinthetext,includingdownhere.The outputfootnoteswilllookexactlylikefootnotesproducedwithmos. †Inthefootnote,thedaggerwillappearwherethefootnotenumberwouldotherwiseappear,asontheleft.

Page2 TheNew-msMacros whichissuppressedexceptinthesismode.Ifyouwantromannumeralpagenumbering,usean ‘‘.afPNi’’request. Thereisanewmacroespeciallyforbibliographyentries,called.XP,whichstandsforexdentedparagraph.Itwillexdentthefirstlineoftheparagraphby\n(PIunits,usually5n(thesameastheindentfor thefirstlineofa.PP).Mostbibliographiesareprintedthisway.Herearesomeexamplesofexdented paragraphs: Lumley,LyleS.,SexinCrustaceans:ShellFishHabits,HarbingerPress,TampaBayandSanDiego, October1979.243pages.Thepioneeringworkinthisfield. Leffadinger,HarryA.,‘‘MolluskMatingSeason:52Weeks,orAllYear?’’inActaBiologica,vol.42,no. 11,November1980.Aprovocativethesis,buttheconclusionsarewrong. Ofcourse,youwillhavetotakecareofitalicizingthebooktitleandjournal,andquotingthetitleofthe journalarticle.IndentationorexdentationcanbechangedbysettingthevalueofnumberregisterPI. Ifyouneedtoproduceendnotesratherthanfootnotes,putthereferencesinafileoftheirown.This issimilartowhatyouwoulddoifyouweretypingthepaperonaconventionaltypewriter.Notethatyou canuseautomaticfootnotenumberingwithoutactuallyhaving.FSand.FEpairsinyourtext.Ifyou placefootnotesinaseparatefile,youcanuse.IPmacroswith\∗∗asahangingtag;thiswillgiveyou numbersattheleft-handmargin.Withsomestylesofendnotes,youwouldwanttouse.PPratherthen .IPmacros,andspecify\∗∗beforethereferencebegins. Therearefournewmacrostohelpproduceatableofcontents.Tableofcontentsentriesmustbe enclosedin.XSand.XEpairs,withoptional.XAmacrosforadditionalentries;argumentsto.XSand.XA specifythepagenumber,tobeprintedattheright.Afinal.PXmacroprintsoutthetableofcontents. Hereisasampleoftypicalinputandoutputtext: .XSii Introduction .XA1 Chapter1:ReviewoftheLiterature .XA23 Chapter2:ExperimentalEvidence .XE .PX TableofContents Introduction...........................................................................................................ii Chapter1:ReviewoftheLiterature........................................................................1 Chapter2:ExperimentalEvidence..........................................................................23 The.XSand.XEpairsmayalsobeusedinthetext,afterasectionheaderforinstance,inwhichcase pagenumbersaresuppliedautomatically.However,mostdocumentsthatrequireatableofcontentsare toolongtoproduceinonerun,whichisnecessaryifthismethodistowork.Itisrecommendedthatyou doatableofcontentsafterfinishingyourdocument.Toprintoutthetableofcontents,usethe.PX macro;ifyouforgetit,nothingwillhappen. Asanaidinproducingtextthatwillformatcorrectlywithbothnroffandtroff,therearesomenew stringdefinitionsthatdefinequotationmarksanddashesforeachofthesetwoformattingprograms.The \∗stringwillyieldtwohyphensinnroff,butintroffitwillproduceanemdash—likethisone.The \∗Qand\∗Ustringswillproduce‘‘and’’introff,but"innroff.(Intypesetting,thedoublequoteis traditionallyconsideredbadform.) Therearenowalargenumberofoptionalforeignaccentmarksdefinedbythemsmacros.Allthe accentmarksavailableinmosarepresent,andtheyallworkjustastheyalwaysdid.However,thereare betterdefinitionsavailablebyplacing.AMatthebeginningofyourdocument.Unlikethemosaccent TheNew-msMacros Page3 marks,theaccentstringsshouldcomeaftertheletterbeingaccented.Hereisalistofthediacritical marks,withexamplesofwhattheylooklike. nameofaccentinputoutput acuteaccente\∗´ e´ graveaccente\∗` e` circumflexo\∗ˆ oˆ cedilla c\∗,c¸ tilde n\∗˜ n˜ question \∗?¿ exclamation\∗!¡ umlaut u\∗:u¨ digraphs\∗8ß hacˇ ek c\∗vcˇ macron a\∗a¯ underdots\∗.s. o-slash o\∗/o/ angstroma\∗oa˚ yogh kni\∗3tkni3t Thorn \∗(ThID thorn \∗(thlp Eth \∗(D--D eth \∗(d--∂ hookedo\∗qoc aeligature\∗(aeæ AEligature\∗(AeÆ oeligature\∗(oeœ OEligature\∗(OeŒ Ifyouwanttousethesenewdiacriticalmarks,don’tforgetthe.AMatthetopofyourfile.Withoutit, somewillnotprintatall,andotherswillbeplacedonthewrongletter. Itisalsopossibletoproducecustomheadersandfootersthataredifferentonevenandoddpages. The.OHand.EHmacrosdefineoddandevenheaders,while.OFand.EFdefineoddandevenfooters. Argumentstothesefourmacrosarespecifiedaswith.tl.Thisdocumentwasproducedwith: .OH´ \fIThe-mxMacros´ ´ Page%\fP´ .EH´ \fIPage%´ ´ The-mxMacros\fP´ Notethatitwouldbeaerrortohaveanapostropheintheheadertext;ifyouneedone,youwillhaveto useadifferentdelimiteraroundtheleft,center,andrightportionsofthetitle.Youcanuseanycharacter asadelimiter,provideditdoesn’tappearelsewhereintheargumentto.OH,.EH,.OF,orEF. Themsmacrosworkinconjunctionwiththetbl,eqn,andreferpreprocessors.Macrostodeal withtheseitemsarereadinonlyasneeded,asarethethesismacros(.TM),thespecialaccentmarkdefinitions(.AM),tableofcontentsmacros(.XSand.XE),andmacrostoformattheoptionalcoverpage. Thecodeforthemspackagelivesin/usr/lib/tmac/tmac.s,andsourcedfilesresideinthedirectory /usr/ucb/lib/ms. 8April1993